Evaluate each of the following if a=3 and b=4.
1)(ab)2
2)ab2
3)(a-b)3

Answer by Abbey(339)

1. (3*4)*2 = 24
2. 3*4*2 = 24
3. (3-4)*3
(-1)*3 = -3